OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution? A. .5 L B. 5 L C. .4 L D. 4 L

OpenStudy (anonymous):

How many liters of alcohol did you start with?

OpenStudy (anonymous):

(hint: 40% of 12)

OpenStudy (anonymous):

I'm confused...

OpenStudy (anonymous):

The concentrations, 40%, 30% are fractions of alcohol-over-water.

OpenStudy (anonymous):

40% alcohol means it is 40 parts alcohol to 100 parts total solution of alcohol+water. So if the concentration is 40% then 40% of the 12 liters is the liters of alcohol in the solution.

OpenStudy (anonymous):

^^ *(sorry, should say, "The concentrations, 40%, 30% are fractions of alcohol volume -over- total volume (alcohol+water).")

OpenStudy (anonymous):

You're starting with 40%, 40% = 4/10, so \[\large \frac{4}{10}=\frac{liters \space of \space alcohol}{total \space volume}\]

OpenStudy (anonymous):

0.4?

OpenStudy (anonymous):

You want 30%, 30% = 3/10, so \[\large \frac{3}{10}=\frac{liters \space of \space alcohol}{12+x \space liters}\]

OpenStudy (anonymous):

Yes, 0.4 means the same as 40%, which means the same as 4/10, you can express that fraction however you want. So, if the total volume you start with is 12 liters, how many of those liters are alcohol?